# Universal Euler-Cartan Circuits for Quantum Field Theories

### v1

@@ Introduction.
Figure 1: (a) Schematic of the proposed algorithm initialized with a state $|{\rm\psi_{i}}\rangle$. The quantum processor (maroon dotted line) is involved in the application of $N$ layers of parametrized unitary operators yielding the state $|{\rm\psi_{f}}\rangle$, followed by measurements of a cost-function and relevant gradients. A classical processor (blue dotted line) verifies the iteration convergence criterion. This checks if the results for the cost-function are converged when compared to that obtained in the previous iteration step for a given $N$. If the iteration convergence criterion is not met, the parameters of the $N$ unitary operators are updated using an optimization routine. Some typical examples are BFGS and the quantum natural gradient (QNG). The updated set of parameters are used in the next iteration and this process is repeated until the desired iteration convergence criterion is reached. Subsequently, a second layer convergence criterion is checked by comparing the result to that obtained with $N-1$ layers. If the convergence criterion is satisfied, the entire process terminates, else the number of layers is grown by 1 and the previous steps are repeated. (b) Decomposition of each layer unitary operator in terms of general two-qubit unitary operators (orange rectangles) and those after Euler and Cartan’s KAK decompositions and compression (orange ellipses). (c) Cartan’s KAK decomposition of a general SU(4) two-qubit unitary operator in terms of four single-qubit SU(2) operators and the SU(4) entangler. (d) Euler’s decomposition of a general SU(2) operator in terms of rotations by three angles about two non-parallel axes (chosen here to be $\hat{y}$ and $\hat{z}$). The SU(4) entangler is decomposed into three controlled-NOT gates and five single-qubit rotations. (e) The two-qubit operator after Euler’s and Cartan’s decompositions used for the optimization process. (f) Legends for the different symbols used. The symbol $R_{\alpha}$ stands for the SU(2) operator $e^{-i\theta\alpha}$, where $\alpha$ is one of the three Pauli matrices and $\theta$ is an unspecified angle determined by the optimization procedure. Two boxes of the same color perform different amounts of rotations about the same axis.
@@
After the application of the $N$ layers, measurements are performed to determine the relevant gradients and the cost-function ${\cal L}(\vec{\theta})$, where $\vec{\theta}$ contains all parameters of the $N$ layer quantum circuit. To obtain the ground state of a given Hamiltonian $H$, the cost-function is chosen to be ${\cal L}(\vec{\theta})=\langle\psi_{\rm f}(\vec{\theta})|H|\psi_{\rm f}(\vec{\theta})\rangle$. To obtain the $n^{\rm th}$ excited state, the cost-function is augmented to ${\cal L}(\vec{\theta})\rightarrow{\cal L}(\vec{\theta})+\sum_{m<n}\lambda_{m}|\langle\psi_{m}|\psi_{\rm f}\rangle|^{2}$, with $\lambda_{m}>0,m=0,\ldots,n-1$. This choice amounts to simultaneously minimizing the energy and the overlap to the $n-1$ lower-energy states that have been already obtained.
@@
$\vec{\theta}_{t+1}=\vec{\theta}_{t}-\eta g^{-1}\frac{\partial{\cal L}}{\partial\vec{\theta}_{t}},$ (1)
where $\eta$ is the learning rate (typically a real number $<1$) and $g$ is the Fubini-Study metric [15, 31]. The latter is given by the real part of the quantum geometric tensor $G$:
@@
Note that both the gradients need to be evaluated using measurements for both the cost-function and the Fubini-Study metric [34, 35]. Subsequently, the updated set of parameters are fed back to the quantum processor and the process is repeated until the iteration convergence criterion is satisfied. Once the latter is satisfied, a second layer convergence criterion is verified. This ensures that the number of layers chosen is sufficient to attain the desired degree of convergence. If this condition is not satisfied, the number of layers is grown by 1 and the earlier steps are repeated. Else, the algorithm terminates. When the number of layers is increased from $N$ to $N+1$, the optimized parameters obtained for the $N$-layer circuit, augmented by an initial guess for the added layer, were used as the choice for the optimization process for $N+1$ layers. This improved the stability of the algorithm.
@@
For $N$ layers and a ring of $L$ qubits, the total number of parameters determined by the above algorithm is $10LN$, where the factor 10 originates from the number of free parameters in the two-qubit unitary operator of Fig. 1(e). Despite being tractable [see Fig. 3(c) for demonstration], conservation of symmetries present in the target Hamiltonian can drastically reduce the number of parameters to be determined. Symmetries can be conserved by modifying the general circuit ansatz of Fig. 1 and allowing only operators that respect this symmetry. Alternatively, the cost-function can be augmented to ensure that the obtained state lies in the correct symmetry sector. For example, states which are invariant under translation by $r$ lattice site(s) can be obtained by starting with a state and circuit ansatz respecting this symmetry. For $r=1(2)$, the number of undetermined parameters for $N$ layers reduces to $10N~{}(20N)$. Alternatively, in order to obtain a ±$\pm 1$ eigenstate of the translation operator $T$, the cost-function can be augmented as: ${\cal L}(\vec{\theta})\rightarrow{\cal L}(\vec{\theta})\mp\mu\langle\psi_{\rm f}(\vec{\theta})|T|\psi_{\rm f}(\vec{\theta})\rangle$, where $\mu$ is a positive real number. Similar considerations apply to symmetry operators related to parity/particle number conservation.
@@
Figure 2: (a) - (c) Results obtained for the ground state energies ($E_{0}$) of the Ising, 3-state Potts and the massive Schwinger models using the optimized Euler-Cartan circuits. A periodic chain of $L$ qubits is subject to $N$ layers for all three models. A translation-invariant Euler-Cartan circuit ansatz is optimized for panels (a, b), while the most general ansatz is considered in panel (c). The different values of $N~{}(L)$ are distinguished by the different colors (markers). For the lowest value of $N$, a uniform choice of $\theta_{0}=0.1$ was used as a starting guess for all of the angles being optimized. Once the iteration convergence criterion is met for a chosen value of $N$, the number of layers is increased by one: $N\rightarrow N+1$. While optimizing for $N+1$ layers, the previously obtained results for the $N$ layers was used as a starting guess, augmented by an initial guess of $\theta_{0}/10$ for the last layer. This process was continued until the layer-convergence criterion was reached. Both convergence criteria were chosen to be $5\times 10^{-4}$. The learning rates were chosen to be $0.04$ for panels (a), (b) and 0.1 for panel (c). (d) Energies for the smallest and largest values of $N$ for $L=16$ are compared to the DMRG results. The smallest choice of $N=2$ layers yielded results within $1\%$ error of the target energies, while the additional layers caused changes in the third and fourth significant digits.
@@ Results.
The proposed algorithm is benchmarked using three paradigmatic one-dimensional QFTs. The results were obtained by modifying standard time-evolution routines for matrix product states. In all three cases, the quantum processor considered comprises a periodic ring with a fixed number of qubits with nearest-neighbor connectivity. This is independent of the model under consideration. The results obtained using the optimized quantum circuit are compared with those from density matrix renormalization group (DMRG) and exact computations.
@@
where $g$ and $h$ were chosen to be $1$ and $\simeq 0.156$ respectively. Figs. 2(a) and 3(a) together present the results for the lowest eight translation-invariant states obtained using the proposed algorithm. The circuit ansatz was chosen to be invariant under translation by 2 sites. Furthermore, the cost-function was augmented to ensure that only $+1$ eigenstates of the operator for single-site translation were obtained. Fig. 2(a) shows the results for the ground state energy for system-sizes between $L=8$ and 20 in steps of 4. The number of layers required to obtain a reasonable approximation of the ground state do not scale with the system-size, which is compatible with the observations of Ref. [31]. In fact, $N=2$ is sufficient to obtain the ground state with error in the energy $<1\%$ [Fig. 2(d)]. Fig. 3(a) shows the results for the first seven excited states for $L=8$. For the chosen convergence criterion of $10^{-3}$, the maximum error in the obtained energies was $0.58\%$. In contrast to the ground state, the number of layers required to achieve this precision ranged between $L$ and $3L$. The different translation-invariant excited states can be classified as lattice counterparts of the single and multiple meson states existing in the QFT. The mass ratio for the lowest two mesonic excitations [square and diamond markers in Fig. 3(a)] obtained for L $=$ 8 obtained using the proposed algorithm is 1.39 which is close to the expected value of 1.41 obtained using exact diagonalization of the corresponding spin-chain. This should be compared with the QFT prediction [36] given by $2\cos(\pi/5)\simeq 1.618$.
@@
Figure 3: (a) - (b) Results for the first seven excited states obtained using optimized Euler-Cartan circuits for the Ising [(a)] and Potts [(b)] models encoded in a periodic chain of $L=8$ qubits. The error percentages when compared to exact diagonalization results (black dotted lines) are shown in the top left insets. The variation of the obtained energies with number of layers is shown in the bottom right insets. The energies were obtained with a convergence criterion of $10^{-3}(10^{-4})$ for the Ising (Potts) models with the same learning rate of 0.02. In both cases, a translation-invariant circuit ansatz was used. The Euler-Cartan circuit for the $i^{\rm th}$ state was obtained by minimizing the Hamiltonian cost-function while orthogonalizing against the previously obtained states. Furthermore, in the Potts model, another term was added to the cost-function to obtain states with $+1$ eigenvalue for the operator $C$ (see maintext). The seventh translation-invariant excited state in the Ising (Potts) models corresponds to the $48^{\rm th}(42^{\rm nd})$ excited state in the spectrum. (c) - (d) Schematic of the confinement phenomena in the Ising and Potts models in the presence of a magnetic field. The latter breaks the two (three)-fold degeneracy of the ground state leading to one true ground state and one (two) false vacua in the Ising (Potts) case. As a consequence, the domain-walls experience a confining force (yellow lines with stripes), leading to the formation of mesonic and baryonic excitations.
@@
$H_{\rm P}=-\sum_{j=1}^{L^{\prime}}\left(\sigma_{j}\sigma_{j+1}^{\dagger}+g\tau+h\sigma+{\rm h.c.}\right),$ (5)
where $\sigma$ and $\tau$ are the shift and clock matrices respectively [44]. Periodic boundary condition was imposed and $L^{\prime}=L/2$ is the number of Potts spins encoded in $L$ qubits. For $g<1,h=0$, the model exhibits a ferromagnetic phase with three degenerate vacua and six domain walls [the three shown in Fig. 3(d) and those obtained by reflection about the domain wall]. The presence of the magnetic field breaks the degeneracy of the ground space, giving rise to one true vacuum and two false vacua. The domain-walls experience a confining potential as in the Ising case. However, now there are not only mesonic excitations formed by two domain walls, but also baryonic excitations formed by three domain walls [Fig. 3(d)]. Notice that $H_{\rm P}$ commutes with the $\mathbb{Z}_{2}$ charge operator $C$ which acts trivially on the state $|0\rangle$ and interchanges $|1\rangle$ and $|2\rangle$. Figs. 2,3 together show the results for $g=h=0.1$ for the lowest eight translation-invariant states with eigenvalue $+1$ for the operator $C$. The 2-site translation-invariant circuit ansatz was chosen with the $C$-eigenvalue imposed using the cost-function. As in the Ising case, the ground state is reached with depth independent of the system-size, with errors in energy less than $0.1\%$ for two layers [Fig. 2(b, d)]. The excited states were obtained using between $L$ and $3L$ layers with maximum error in the obtained energies being $2.26\%$ [Fig. 3(b)]. The first excited state is one of the false vacua, with the rest comprising mesonic and baryonic excitations. This is verified by counting the average number of domain-walls in the corresponding state.
@@
where $P_{j}=[1+(-1)^{j}Z_{j}]/2$. In contrast to the Ising and Potts models, this model involves an all-to-all coupling Hamiltonian. To obtain the ground state of this model, the general circuit ansatz of Fig. 1 is used. Note the existence of the conserved U(1) charge $Q$$[H,Q]=0$, where $Q=\sum_{j=1}^{L}Z_{j}$. To test the algorithm under the most challenging of conditions, the ground states were obtained for system sizes between L = 8 and 16 without conserving this charge. For these simulations, $m=0.5,g=0.3$ were chosen. Despite the long-range nature of the model Hamiltonian, reasonable approximations of the ground-state were obtained with circuit-depths that are independent of the system-size with energies within $1\%$ accuracy obtained within application of 2 layers. Excited states were also obtained similar to the Ising and Potts models and are not shown here for brevity.

### v2

@@ Introduction.
Figure 1: (a) Schematic of the proposed algorithm initialized with a state $|{\rm\psi_{i}}\rangle$. The quantum processor (maroon dotted line) is involved in the application of $N$ layers of parametrized unitary operators yielding the state $|{\rm\psi_{f}}\rangle$, followed by measurements of a cost-function and relevant gradients. A classical processor (blue dotted line) verifies the iteration convergence criterion. This checks if the results for the cost-function are converged when compared to that obtained in the previous iteration step for a given NN. If the iteration convergence criterion is not met, the parameters of the $N$ unitary operators are updated using an optimization routine. Some typical examples are BFGS and the quantum natural gradient (QNG). The updated set of parameters are used in the next iteration and this process is repeated until the desired iteration convergence criterion is reached. Subsequently, a second layer convergence criterion is checked by comparing the result to that obtained with N1N-1 layers. If the convergence criterion is satisfied, the entire process terminates, else the number of layers is grown by 1 and the previous steps are repeated. (b) Decomposition of each layer unitary operator in terms of general two-qubit unitary operators (orange rectangles) and those after Euler and Cartan’s KAK decompositions and compression (orange ellipses). (c) Cartan’s KAK decomposition of a general SU(4) two-qubit unitary operator in terms of four single-qubit SU(2) operators and the SU(4) entangler. (d) Euler’s decomposition of a general SU(2) operator in terms of rotations by three angles about two non-parallel axes (chosen here to be $\hat{y}$ and $\hat{z}$). The SU(4) entangler is decomposed into three controlled-NOT gates and five single-qubit rotations. (e) The two-qubit operator after Euler’s and Cartan’s decompositions used for the optimization process. (f) Legends for the different symbols used. The symbol $R_{\alpha}$ stands for the SU(2) operator $e^{-i\theta\alpha}$, where $\alpha$ is one of the three Pauli matrices and $\theta$ is an unspecified angle determined by the optimization procedure. Two boxes of the same color perform different amounts of rotations about the same axis.
@@
After the application of the $N$ layers, measurements are performed to determine the relevant gradients and the cost-function ${\cal L}(\vec{\theta})$, where $\vec{\theta}$ contains all parameters of the $N$ layer quantum circuit. To obtain the ground state of a given Hamiltonian $H$, the cost-function is chosen to be ${\cal L}(\vec{\theta})=\langle\psi_{\rm f}(\vec{\theta})|H|\psi_{\rm f}(\vec{\theta})\rangle$. To obtain the $n^{\rm th}$ excited state, the cost-function is augmented to ${\cal L}(\vec{\theta})\rightarrow{\cal L}(\vec{\theta})+\sum_{m<n}\lambda_{m}|\langle\psi_{m}|\psi_{\rm f}\rangle|^{2}$, with $\lambda_{m}>0,m=0,\ldots,n-1$. This choice amounts to simultaneously minimizing the energy and the overlap to the n1n-1 lower-energy states that have been already obtained.
@@
$\vec{\theta}_{t+1}=\vec{\theta}_{t}-\eta g_{\rm FS}^{-1}\frac{\partial{\cal L}}{\partial\vec{\theta}_{t}},$ (1)
where $\eta$ is the learning rate (typically a real number <1<1) and $g_{\rm FS}$ is the Fubini-Study metric [15, 31]. The latter is given by the real part of the quantum geometric tensor $G$:
@@
Note that both the gradients need to be evaluated using measurements for both the cost-function and the Fubini-Study metric [34]. Subsequently, the updated set of parameters are fed back to the quantum processor and the process is repeated until the iteration convergence criterion is satisfied. Once the latter is satisfied, a second layer convergence criterion is verified. This ensures that the number of layers chosen is sufficient to attain the desired degree of convergence. If this condition is not satisfied, the number of layers is grown by 1 and the earlier steps are repeated. Else, the algorithm terminates. When the number of layers is increased from $N$ to $N+1$, the optimized parameters obtained for the $N$-layer circuit, augmented by an initial guess for the added layer, were used as the choice for the optimization process for $N+1$ layers. This improved the stability of the algorithm.
@@
For $N$ layers and a ring of $L$ qubits, the total number of parameters determined by the above algorithm is $10LN$, where the factor 10 originates from the number of free parameters in the two-qubit unitary operator of Fig. 1(e). The number of entangling CNOT gates per qubit is given by $3N$. Despite being tractable [see Fig. 3(c) for demonstration], conservation of symmetries present in the target Hamiltonian can drastically reduce the number of parameters to be determined. Symmetries can be conserved by modifying the general circuit ansatz of Fig. 1 and allowing only operators that respect this symmetry. Alternatively, the cost-function can be augmented to ensure that the obtained state lies in the correct symmetry sector. For example, states which are invariant under translation by $r$ lattice site(s) can be obtained by starting with a state and circuit ansatz respecting this symmetry. For $r=1(2)$, the number of undetermined parameters for $N$ layers reduces to 10N(20N)10N~(20N). Alternatively, in order to obtain a $\pm 1$ eigenstate of the translation operator $T$, the cost-function can be augmented as: ${\cal L}(\vec{\theta})\rightarrow{\cal L}(\vec{\theta})\mp\mu\langle\psi_{\rm f}(\vec{\theta})|T|\psi_{\rm f}(\vec{\theta})\rangle$, where $\mu$ is a positive real number. Similar considerations apply to symmetry operators related to parity/particle number conservation.
Note that the described circuit ansatz can be modified to incorporate constraints arising from specific hardware architectures. For instance, if all-to-all coupling is permitted (applicable, for instance, to trapped-ion based quantum computing platforms), the proposed ansatz can be straightforwardly generalized to take advantage of the additional couplings. In this work, the simplest case of only nearest-neighbor coupling is considered.
@@
Figure 2: (a) - (c) Results obtained for the ground state energies ($E_{0}$) of the Ising, 3-state Potts and the massive Schwinger models using the optimized Euler-Cartan circuits. A periodic chain of $L$ qubits is subject to NN layers for all three models. A translation-invariant Euler-Cartan circuit ansatz is optimized for panels (a, b), while the most general ansatz is considered in panel (c). The different values of N(L)N~(L) are distinguished by the different colors (markers). For the lowest value of $N$, a uniform choice of $\theta_{0}=0.1$ was used as a starting guess for all of the angles being optimized. Once the iteration convergence criterion is met for a chosen value of $N$, the number of layers is increased by one: $N\rightarrow N+1$. While optimizing for $N+1$ layers, the previously obtained results for the $N$ layers was used as a starting guess, augmented by an initial guess of $\theta_{0}/10$ for the last layer. This process was continued until the layer-convergence criterion was reached. Both convergence criteria were chosen to be $5\times 10^{-4}$. The learning rates were chosen to be $0.04$ for panels (a), (b) and 0.1 for panel (c). (d) Energies and fidelities to the corresponding target states for the smallest and largest values of NN for $L=16$ are compared to the DMRG results. The smallest choice of $N=2$ layers yielded results within $1\%$ error of the target energies, while the additional layers caused changes in the third and fourth significant digits. While the precision for the massive Schwinger model is lower due to the all-to-all couplings in the spin Hamiltonian [Eq. (6)] when compared to the other two models, increasing the number of layers leads to further improvement in the obtained results (see also Ref. [35]). Note that the exact values of the changes in the energies with layers depend on the specific choices of the coupling constants and the distances between the initial and target ground states.
@@ Results.
The proposed algorithm is benchmarked using three paradigmatic one-dimensional QFTs. The results were obtained by modifying standard time-evolution routines for matrix product states. In all three cases, the quantum processor considered comprises a periodic ring with a fixed number of qubits with nearest-neighbor connectivity. This is independent of the model under consideration. The results obtained using the optimized quantum circuit are compared with those from density matrix renormalization group (DMRG) and exact computations. While the results shown below make use of the QNG optimizer, the latter is not essential for the proposed circuit ansatz. For instance, both Ref. [35] and Ref. [36] use ADAM [14] or its variant for unitary manifolds for the current ansatz.
@@
where $g$ and $h$ were chosen to be $1$ and $\simeq 0.154$ respectively. Figs. 2(a) and 3(a) together present the results for the lowest eight translation-invariant states obtained using the proposed algorithm. The circuit ansatz was chosen to be invariant under translation by 2 sites. Furthermore, the cost-function was augmented to ensure that only +1+1 eigenstates of the operator for single-site translation were obtained. Fig. 2(a) shows the results for the ground state energy for system-sizes between $L=8$ and 20 in steps of 4. The number of layers required to obtain a reasonable approximation of the ground state do not scale with the system-size. This is compatible with the observations of Ref. [31] and is a consequence of the ‘relatively-low amount of entanglement’ in ground states of one-dimensional models with gapped spectrum [41]. In fact, $N=2$ is sufficient to obtain the ground state with error in the energy $1\%$ [Fig. 2(d)]. Fig. 3(a) shows the results for the first seven excited states for $L=8$. For the chosen convergence criterion of $10^{-3}$, the maximum error in the obtained energies was $0.58\%$. In contrast to the ground state, the number of layers required to achieve this precision ranged between $L$ and $3L$. The different translation-invariant excited states can be classified as lattice counterparts of the single and multiple meson states existing in the QFT. The mass ratio for the lowest two mesonic excitations [square and diamond markers in Fig. 3(a)] obtained for L == 8 obtained using the proposed algorithm is 1.39 which is close to the expected value of 1.41 obtained using exact diagonalization of the corresponding spin-chain. This should be compared with the QFT prediction [37] given by $2\cos(\pi/5)\simeq 1.618$.
@@
Figure 3: (a) - (b) Results for the first seven excited states obtained using optimized Euler-Cartan circuits for the Ising [(a)] and Potts [(b)] models encoded in a periodic chain of $L=8$ qubits. The error percentages when compared to exact diagonalization results (black dotted lines) are shown in the top left insets. The variation of the obtained energies with number of layers is shown in the bottom right insets. The energies were obtained with a convergence criterion of $10^{-3}(10^{-4})$ for the Ising (Potts) models with the same learning rate of 0.02. In both cases, a translation-invariant circuit ansatz was used. The Euler-Cartan circuit for the $i^{\rm th}$ state was obtained by minimizing the Hamiltonian cost-function while orthogonalizing against the previously obtained states. Furthermore, in the Potts model, another term was added to the cost-function to obtain states with +1+1 eigenvalue for the operator $C$ (see maintext). The seventh translation-invariant excited state in the Ising (Potts) models corresponds to the $48^{\rm th}(42^{\rm nd})$ excited state in the spectrum. (c) - (d) Schematic of the confinement phenomena in the Ising and Potts models in the presence of a magnetic field. The latter breaks the two (three)-fold degeneracy of the ground state leading to one true ground state and one (two) false vacua in the Ising (Potts) case. As a consequence, the domain-walls experience a confining force (yellow lines with stripes), leading to the formation of mesonic and baryonic excitations. The non-monotonicity in the convergence for highly excited states can be improved by lowering the learning rates used in the circuit optimization process.
@@
$H_{\rm P}=-\frac{2}{3\sqrt{3}}\sum_{j=1}^{L^{\prime}}\left(\sigma_{j}\sigma_{j+1}^{\dagger}+g\tau_{j}+h\sigma_{j}+{\rm h.c.}\right),$ (5)
where $\sigma$ and $\tau$ are the shift and clock matrices respectively [46]. Periodic boundary condition was imposed and $L^{\prime}=L/2$ is the number of Potts spins encoded in $L$ qubits. For $g<1,h=0$, the model exhibits a ferromagnetic phase with three degenerate vacua and six domain walls [the three shown in Fig. 3(d) and those obtained by reflection about the domain wall]. The presence of the magnetic field breaks the degeneracy of the ground space, giving rise to one true vacuum and two false vacua. The domain-walls experience a confining potential as in the Ising case. However, now there are not only mesonic excitations formed by two domain walls, but also baryonic excitations formed by three domain walls [Fig. 3(d)]. Notice that $H_{\rm P}$ commutes with the $\mathbb{Z}_{2}$ charge operator $C$ which acts trivially on the state $|0\rangle$ and interchanges $|1\rangle$ and $|2\rangle$. Figs. 2,3 together show the results for $g=h=0.1$ for the lowest eight translation-invariant states with eigenvalue +1+1 for the operator $C$. The 2-site translation-invariant circuit ansatz was chosen with the $C$-eigenvalue imposed using the cost-function. As in the Ising case, the ground state is reached with depth independent of the system-size, with errors in energy less than $0.1\%$ for two layers [Fig. 2(b, d)]. The excited states were obtained using between $L$ and $3L$ layers with maximum error in the obtained energies being $2.26\%$ [Fig. 3(b)]. The first excited state is one of the false vacua, with the rest comprising mesonic and baryonic excitations. This is verified by counting the average number of domain-walls in the corresponding state.
@@
where $P_{j}=[1+(-1)^{j}Z_{j}]/2$. In contrast to the Ising and Potts models, this model involves an all-to-all coupling Hamiltonian. To obtain the ground state of this model, the general circuit ansatz of Fig. 1 is used. Note the existence of the conserved U(1) charge QQ[H,Q]=0[H,Q]=0, where $Q=\sum_{j=1}^{L}Z_{j}$. To test the algorithm under the most challenging of conditions, the ground states were obtained for system sizes between L = 8 and 16 without conserving this charge. For these simulations, $m=0.5,g=0.3$ were chosen. Despite the long-range nature of the model Hamiltonian, reasonable approximations of the ground-state were obtained with circuit-depths that are independent of the system-size with energies within $1\%$ accuracy obtained within application of 2 layers. Excited states were also obtained similar to the Ising and Potts models and are not shown here for brevity. In the three examples analyzed above, the number of iterations required for convergence for a given number of layers is ${\cal O}(10^{2})$.